# From Reusing to Forecasting: Accelerating Diffusion Models with TaylorSeers

### v1

@@ 2.2 Denoising Network Acceleration
Feature caching has emerged as particularly advantageous due to its training-free implementation. Initially developed for U-Net architectures [17, 26], caching mechanisms have evolved to address the computational demands of DiT models, which offer superior performance but at increased computational cost. Advanced techniques such as FORA [37] and -DiT [5] leverage attention and MLP representation reuse, while TeaCache [21] implements dynamic timestep-dependent difference estimation to optimize caching decisions. DiTFastAttn [46] reduces redundancies in self-attention computation across spatial, temporal, and conditional dimensions. The ToCa series [53, 54] mitigates information loss through dynamic feature updates, while EOC [30] introduces an error-optimized framework utilizing prior knowledge extraction and adaptive optimization. Recent innovations including UniCP [42], which integrates dynamic cache window adjustment with pruning, and RAS [23], which implements region-adaptive sampling rates based on model focus, further enhance computational efficiency while preserving generation fidelity.
@@ 3.2 TaylorSeer
Unlike Naïve Feature Caching, TaylorSeer transforms simply feature reuse into a predictive process, explicitly modeling feature trajectories through high-order finite differences. This key transition from directly caching to forecasting not only significantly improves prediction accuracy but also enables efficient long-range inference.. As a result, TaylorSeer provides a robust and scalable solution for accelerating diffusion model inference, maintaining high generation quality even at large acceleration ratios.
@@ 4 Experiments
| Method | Efficient | Acceleration | Image Reward | CLIP | |||
| FLUX.1[16] | Attention [7] | Latency(s) | Speed | FLOPs(T) | Speed | DrawBench | Score |
| [dev]: 50 steps | ✔ | 17.20 | 1.00 | 3719.50 | 1.00 | 0.9898 | 19.604 |
| steps | ✔ | 10.49 | 1.64 | 2231.70 | 1.67 | 0.9739 | 19.526 |
| -DiT () † | ✔ | 11.87 | 1.45 | 2480.01 | 1.50 | 0.9316 | 19.350 |
| steps † | ✔ | 8.80 | 1.95 | 1859.75 | 2.00 | 0.9429 | 19.325 |
| steps † | ✔ | 7.11 | 2.42 | 1487.80 | 2.62 | 0.9317 | 19.027 |
| steps † | ✔ | 6.09 | 2.82 | 1264.63 | 3.13 | 0.9346 | 18.904 |
| -DiT () † | ✔ | 8.81 | 1686.76 | 2.21 | 0.8561 | 18.833 | |
| FORA † [37] | ✔ | 7.08 | 2.43 | 1320.07 | 2.82 | 0.9227 | 18.950 |
| ToCa [53] | ✘ | 10.80 | 1.59 | 1126.76 | 3.30 | 0.9731 | 19.030 |
| [54] | ✔ | 5.88 | 2.93 | 1078.34 | 3.45 | 0.9896 | 19.595 |
| TaylorSeer | ✔ | 5.82 | 2.96 | 893.54 | 4.16 | 1.0296 | 19.437 |
| FORA † [37] | ✔ | 5.43 | 3.17 | 967.91 | 3.84 | 0.8675 | 18.560 |
| ToCa † [53] | ✘ | 8.47 | 2.03 | 784.54 | 4.74 | 0.9086 | 18.380 |
| DuCa † [54] | ✔ | 4.89 | 3.52 | 816.55 | 4.56 | 0.9470 | 19.082 |
| TaylorSeer | ✔ | 4.87 | 3.53 | 744.81 | 4.99 | 0.9953 | 19.637 |
| TaylorSeer | ✔ | 5.19 | 3.31 | 744.81 | 4.99 | 1.0039 | 19.427 |
@@
| Method | Efficient | Acceleration | VBench | |||
| HunyuanVideo[43] | Attention [7] | Latency(s) | Speed | FLOPs(T) | Speed | Score(%) |
| Original: 50 steps | ✔ | 318.24 | 1.00 | 29773.0 | 1.00 | 80.66 |
| steps | ✔ | 70.34 | 4.52 | 6550.1 | 4.55 | 78.74 |
| FORA [37] | ✔ | 67.19 | 4.74 | 5960.4 | 5.00 | 78.83 |
| ToCa [53] | ✘ | 77.82* | 4.09 | 7006.2 | 4.25 | 78.86 |
| DuCa [54] | ✔ | 71.10 | 4.48 | 6483.2 | 4.62 | 78.72 |
| TaylorSeer | ✔ | 68.42 | 4.65 | 5960.4 | 5.00 | 79.93 |
| TaylorSeer | ✔ | 61.99 | 5.13 | 5359.1 | 5.56 | 79.78 |
@@ 4.1 Experiment Settings
For the text-to-image generation task, we perform inference on 200 DrawBench [35] prompts to generate images with a resolution of 1000 1000. We then evaluate the generated samples using Image Reward [44] and CLIP Score [9] as key metrics to assess image quality and text alignment. For the text-to-video generation task, we leverage the VBench [13] evaluation framework, utilizing its 946 benchmark prompts. For each prompt, we generate five samples with different random seeds, totaling 4,730 videos. We then systematically assess the generated results based on 16 core evaluation dimensions defined by the VBench framework to provide a comprehensive evaluation of the model’s performance. For the class-conditional image generation task, we uniformly sample from 1,000 ImageNet [33] categories, generating 50,000 images with a resolution of 256 256. We use FID-50k [10] as the primary evaluation metric, complemented by sFID and Inception Score for robust evaluation.
@@ 4.2 Text-to-Image Generation
We compared TaylorSeer with existing methods. While DuCa [54] (=5) achieves a 3.45 FLOPs speedup with an Image Reward of 0.9896, and ToCa [53] (=5) offers 3.30 acceleration with reduced quality (0.9731), TaylorSeer (=6, =1) significantly outperforms them. At 4.99 acceleration, it maintains superior Image Reward (0.9953) and CLIP Score (19.637). Notably, when pushing acceleration further, baseline methods suffer severe quality degradation: DuCa (=6) drops to 0.9470 Image Reward at 4.56 acceleration, while ToCa (=8) plummets to 0.9086 at 4.74 acceleration. In contrast, TaylorSeer (=5, =2) sustains 1.0296 Image Reward even at 4.16 acceleration, demonstrating unmatched efficiency-fidelity balance.
Qualitative Study.
@@ 4.3 Text-to-Video Generation
On HunyuanVideo, TaylorSeer with =5 and =1 reduces inference latency to 68 seconds and computational cost to 5960.4TFLOPs ( speedup), achieving a 79.93% VBench score that outperforms both ToCa and DuCa. With =6 and =1, performance improves further to speedup while maintaining a 79.78% score. These results show that TaylorSeer preserves temporal consistency and reduces computational demands, making high-quality video synthesis more efficient.
@@ 4.4 Class-Conditional Image Generation
We compared TaylorSeer with methods such as ToCa [53], FORA [37], DuCa [54], and reduced DDIM steps on DiT-XL/2 [28], demonstrating that TaylorSeer significantly outperforms the others in both acceleration ratio and generation quality. TaylorSeer(=3, =3) achieves an FID-50k of 2.34 while providing a 2.77 acceleration, showing that higher-order Taylor expansions preserve feature quality at high acceleration ratios without incurring additional computational overhead. At a 4.53 acceleration, our method (=5, =3) maintains an FID of 2.65, outperforming state-of-the-art models such as ToCa [53] and DuCa [54]. Notably, as the acceleration ratio increases beyond 3.5, other methods like FORA, ToCa, and DuCa exhibit significant degradation in FID, leading to severe deterioration in image quality, while TaylorSeer consistently maintains superior performance. This robustness under extreme acceleration conditions further validates the effectiveness of our approach in predicting future features, which minimizes error accumulation compared to cache-then-reuse strategies that struggle to adapt to higher acceleration demands.
@@ 4.5 Ablation Studies
We conduct ablation experiments on DiT-XL/2 [28] to evaluate TaylorSeer, focusing on the impact of the interval parameter and the Taylor expansion order on computational efficiency and generation quality. Results indicate that when =0 (direct feature reuse), performance degrades rapidly as increases. Introducing higher-order Taylor expansions significantly improves generation quality, especially in long-interval scenarios. Even a first-order approximation (=1) substantially enhances performance, reducing FID from non-generation to 4.29 at =7. Higher-order expansions () further improve generation quality, capturing nonlinear feature variations effectively. In high-acceleration settings (=5 or ), third-order expansion maintains low FID (2.78–3.15), outperforming direct reuse. While continues to refine results, improvements saturate beyond the third order, as observed in both sFID and FID trends.
Overall, TaylorSeer demonstrates that Taylor expansion-based forecasting effectively balances efficiency and quality, particularly in high-acceleration scenarios. Higher-order expansions () enable high-quality generation with reduced computational cost, making the method suitable for real-time or resource-constrained applications. Detailed results are provided in the supplementary materials.
@@ 8 Anonymous Page for Video Presentation
To further showcase the advantages of TaylorCache in video generation, we have created an anonymous GitHub page. For a more detailed demonstration, please visit https://taylorseer.github.io/TaylorSeer/. Additionally, the videos are also available in the Supplementary Material.

### v2

@@ 2.2 Denoising Network Acceleration
Feature caching has emerged as particularly advantageous due to its training-free implementation. Initially developed for U-Net architectures [19, 30], caching mechanisms have evolved to address the computational demands of DiT models, which offer superior performance but at increased computational cost. Advanced techniques such as FORA [41] and -DiT [6] leverage attention and MLP representation reuse, while TeaCache [23] implements dynamic timestep-dependent difference estimation to optimize caching decisions. DiTFastAttn [53] reduces redundancies in self-attention computation across spatial, temporal, and conditional dimensions. The ToCa series [64, 63, 24, 61] mitigates information loss through dynamic feature updates, while EOC [34] introduces an error-optimized framework utilizing prior knowledge extraction and adaptive optimization. FasterCache [29] extends caching to video DiTs through dynamic feature reuse and a CFG-Cache strategy. Recent innovations including UniCP [46], which integrates dynamic cache window adjustment with pruning, and RAS [26], which implements region-adaptive sampling rates based on model focus, further enhance computational efficiency while preserving generation fidelity. Caching has also been widely adopted in other tasks such as image editing [51].
@@ 3.2 TaylorSeer
Unlike Naïve Feature Caching, TaylorSeer transforms simply feature reuse into a predictive process, explicitly modeling feature trajectories through high-order finite differences. This key transition from directly caching to forecasting not only significantly improves prediction accuracy but also enables efficient long-range inference. As a result, TaylorSeer provides a robust and scalable solution for accelerating diffusion model inference, maintaining high generation quality even at large acceleration ratios.
@@ 4 Experiments
| Method | Efficient | Acceleration | Image Reward | CLIP | PSNR | SSIM | LPIPS | |||
| FLUX.1[18] | Attention [8] | Latency(s) | Speed | FLOPs(T) | Speed | DrawBench | Score | |||
| [dev]: 50 steps | ✔ | 17.20 | 1.00 | 3719.50 | 1.00 | 0.9898 | 19.604 | - | - | - |
| steps | ✔ | 10.49 | 1.64 | 2231.70 | 1.67 | 0.9739 | 19.526 | 30.310 | 0.7819 | 0.2461 |
| steps † | ✔ | 8.80 | 1.95 | 1859.75 | 2.00 | 0.9429 | 19.325 | 29.576 | 0.7337 | 0.3106 |
| steps † | ✔ | 7.11 | 2.42 | 1487.80 | 2.62 | 0.9317 | 19.027 | 29.122 | 0.6971 | 0.3619 |
| steps † | ✔ | 6.09 | 2.82 | 1264.63 | 3.13 | 0.9346 | 18.904 | 28.881 | 0.6776 | 0.3913 |
| -DiT () † | ✔ | 11.87 | 1.45 | 2480.01 | 1.50 | 0.9316 | 19.350 | 29.399 | 0.7227 | 0.3301 |
| -DiT () † | ✔ | 8.81 | 1686.76 | 2.21 | 0.8561 | 18.833 | 28.794 | 0.6665 | 0.4133 | |
| FORA † [41] | ✔ | 7.08 | 2.43 | 1320.07 | 2.82 | 0.9227 | 18.950 | 30.652 | 0.7666 | 0.2450 |
| TeaCache † [23] | ✔ | 7.56 | 1412.90 | 2.63 | 0.9687 | 19.096 | 28.154 | 0.4674 | 0.6624 | |
| TaylorSeer | ✔ | 6.59 | 2.61 | 1320.07 | 2.82 | 1.0181 | 19.397 | 30.762 | 0.7818 | 0.2300 |
| FORA † [41] | ✔ | 5.17 | 3.33 | 893.54 | 4.16 | 0.8235 | 18.279 | 28.430 | 0.6030 | 0.4915 |
| ToCa [64] | ✘ | 10.80 | 1.59 | 1126.76 | 3.30 | 0.9731 | 19.030 | 29.642 | 0.7082 | 0.3319 |
| [63] | ✔ | 5.88 | 2.93 | 1078.34 | 3.45 | 0.9896 | 19.595 | 29.413 | 0.7142 | 0.3082 |
| TeaCache † [23] | ✔ | 6.07 | 2.83 | 1115.44 | 3.33 | 0.9670 | 18.783 | 28.147 | 0.4674 | 0.6624 |
| TaylorSeer | ✔ | 5.61 | 3.07 | 967.91 | 3.84 | 1.0204 | 19.433 | 29.863 | 0.7273 | 0.2993 |
| TaylorSeer | ✔ | 5.82 | 2.96 | 893.54 | 4.16 | 1.0296 | 19.437 | 29.330 | 0.6994 | 0.3456 |
| FORA † [41] | ✔ | 4.53 | 3.80 | 744.81 | 4.99 | 0.7761 | 17.986 | 28.360 | 0.6001 | 0.5177 |
| ToCa † [64] | ✘ | 8.47 | 2.03 | 784.54 | 4.74 | 0.9086 | 18.380 | 28.961 | 0.6248 | 0.4470 |
| DuCa † [63] | ✔ | 4.89 | 3.52 | 816.55 | 4.56 | 0.9470 | 19.082 | 28.672 | 0.6228 | 0.4182 |
| TeaCache † [23] | ✔ | 4.98 | 3.58 | 892.35 | 4.17 | 0.8683 | 18.500 | 28.146 | 0.4692 | 0.6640 |
| TaylorSeer | ✔ | 4.87 | 3.53 | 744.81 | 4.99 | 0.9953 | 19.637 | 28.826 | 0.6544 | 0.4097 |
| TaylorSeer | ✔ | 5.19 | 3.31 | 744.81 | 4.99 | 1.0039 | 19.427 | 28.945 | 0.6556 | 0.4020 |
@@
| Method | Efficient | Acceleration | VBench | Quality Metrics | |||||
| HunyuanVideo[47] | Attention [8] | Latency(s) | Speed | FLOPs(T) | Speed | Score(%) | PSNR | SSIM | LPIPS |
| Original: 50 steps | ✔ | 318.24 | 1.00 | 29773.0 | 1.00 | 80.66 | - | - | - |
| steps | ✔ | 70.34 | 4.55 | 6550.1 | 4.55 | 78.74 | 16.816 | 0.5708 | 0.3990 |
| AdaCache [15] | ✔ | 120.59 | 2.64 | 11182.5 | 2.66 | 80.25 | 17.089 | 0.6958 | 0.3197 |
| [41] | ✔ | 112.57 | 2.83 | 9929.2 | 3.00 | 80.26 | 19.447 | 0.7688 | 0.1960 |
| TeaCache [23] | ✔ | 109.70 | 2.90 | 10023.9 | 2.97 | 80.44 | 20.230 | 0.7927 | 0.1759 |
| ✔ | 113.44 | 2.81 | 9929.2 | 3.00 | 80.74 | 21.423 | 0.8335 | 0.1353 | |
| [41] | ✔ | 67.19 | 4.74 | 5960.4 | 5.00 | 78.83 | 16.072 | 0.6334 | 0.3457 |
| ToCa [64] | ✘ | 77.82* | 4.09 | 7006.2 | 4.25 | 78.86 | 16.143 | 0.6424 | 0.3432 |
| DuCa [63] | ✔ | 71.10 | 4.58 | 6483.2 | 4.48 | 78.72 | 16.169 | 0.6443 | 0.3516 |
| TeaCache [23] | ✔ | 67.36 | 6550.1 | 4.55 | 79.36 | 16.072 | 0.6216 | 0.4377 | |
| TaylorSeer | ✔ | 68.42 | 4.65 | 5960.4 | 5.00 | 79.93 | 16.796 | 0.7039 | 0.2691 |
| TaylorSeer | ✔ | 61.99 | 5.13 | 5359.1 | 5.56 | 79.78 | 16.241 | 0.7215 | 0.3009 |
@@ 4.1 Experiment Settings
For the text-to-image generation task, we perform inference on 200 DrawBench [39] prompts to generate images with a resolution of 1000 1000. We then evaluate the generated samples using ImageReward [50] and CLIP Score [10] as key metrics to assess image quality and text alignment. For the text-to-video generation task, we leverage the VBench [14] evaluation framework, utilizing its 946 benchmark prompts. For each prompt, we generate five samples with different random seeds, totaling 4,730 videos. We then systematically assess the generated results based on 16 core evaluation dimensions defined by the VBench framework to provide a comprehensive evaluation of the model’s performance. For the class-conditional image generation task, we uniformly sample from 1,000 ImageNet [37] categories, generating 50,000 images with a resolution of 256 256. We use FID-50k [11] as the primary evaluation metric, complemented by sFID and Inception Score for robust evaluation. We assess the fidelity of our accelerated outputs against the original results using PSNR, SSIM [49], and LPIPS [57].
@@ 4.2 Text-to-Image Generation
We compared TaylorSeer with existing methods. While DuCa (=5) achieves a 3.45 speedup with an ImageReward of 0.9896 and ToCa (=5) offers 3.30 acceleration at reduced quality (0.9731), TaylorSeer already demonstrates superior performance in this range. The performance gap widens significantly at higher accelerations. When pushed towards 5 speedup, baseline methods suffer severe quality degradation; for instance, DuCa (=6) and ToCa (=8) see their ImageReward drop to 0.9470 and 0.9086, respectively. In stark contrast, TaylorSeer (=6, =2) sustains an exceptional 1.0039 ImageReward at a 4.99 acceleration, consistently outperforming all competitors in PSNR, SSIM, and LPIPS as well.
@@ 4.3 Text-to-Video Generation
On HunyuanVideo, TaylorSeer with =5 and =1 reduces inference latency to 68 seconds and computational cost to 5960.4TFLOPs ( speedup), achieving a 79.93% VBench score that outperforms both ToCa and DuCa. With =6 and =1, performance improves further to speedup while maintaining a 79.78% score. It also achieves closer-to-original outputs with better PSNR, SSIM, and LPIPS than all baselines.
@@ 4.4 Class-Conditional Image Generation
We compared TaylorSeer with methods such as ToCa, FORA, DuCa, and reduced DDIM steps on DiT-XL/2 [32], demonstrating that TaylorSeer significantly outperforms the others in both acceleration ratio and generation quality. TaylorSeer(=3, =3) achieves an FID-50k of 2.34 while providing a 2.77 acceleration, showing that higher-order Taylor expansions preserve feature quality at high acceleration ratios without incurring additional computational overhead. At a 4.53 acceleration, our method (=5, =3) maintains an FID of 2.65, outperforming state-of-the-art models such as ToCa and DuCa. Notably, as the acceleration ratio increases beyond 3.5, other methods like FORA, ToCa, and DuCa exhibit significant degradation in FID, leading to severe deterioration in image quality, while TaylorSeer consistently maintains superior performance. This robustness under extreme acceleration conditions further validates the effectiveness of our approach in predicting future features, which minimizes error accumulation compared to cache-then-reuse strategies that struggle to adapt to higher acceleration demands.
@@ 4.5 Ablation Studies
We conduct ablation experiments on DiT-XL/2, focusing on the impact of the interval parameter and the Taylor expansion order on computational efficiency and generation quality. Results indicate that when =0 (direct feature reuse), performance degrades rapidly as increases. Introducing higher-order Taylor expansions significantly improves generation quality, especially in long-interval scenarios. Even a first-order approximation (=1) substantially enhances performance, reducing FID from non-generation to 4.29 at =7. Higher-order expansions () further improve generation quality, capturing nonlinear feature variations effectively. In high-acceleration settings (=5 or ), third-order expansion maintains low FID (2.78–3.15), outperforming direct reuse. While continues to refine results, improvements saturate beyond the third order, as observed in both sFID and FID trends. Overall, TaylorSeer demonstrates that Taylor expansion-based forecasting effectively balances efficiency and quality, particularly in high-acceleration scenarios. Higher-order expansions () enable high-quality generation with reduced computational cost, making the method suitable for real-time or resource-constrained applications. Detailed results are provided in the supplementary materials.
@@ 5 Conclusion
Acknowledgement
This work was partially supported by the Dream Set Off - Kunpeng&Ascend Seed Program.
@@ 8 Anonymous Page for Video Presentation
To further showcase the advantages of TaylorSeer in video generation, we have created an anonymous GitHub page. For a more detailed demonstration, please visit https://taylorseer.github.io/TaylorSeer/. Additionally, the videos are also available in the Supplementary Material.
@@ 10 Supplementary Visualization of Feature Trajectories in Diffusion Models
11 More Results for TaylorSeer
To further validate the broad applicability and adaptability of TaylorSeer, we conducted comprehensive experiments on several other mainstream text-to-video DiT models, including HiDream [3], FramePack [56], and WAN2.1 [48]. The results demonstrate that TaylorSeer consistently achieves superior performance across these platforms.
As a representative example, Table 5 presents a detailed performance comparison on the HiDream model. While achieving a 4.0 acceleration, TaylorSeer significantly outperforms the competing method, TeaCache, across all key evaluation metrics, including PSNR, SSIM, LPIPS, and ImageReward.
Furthermore, our project page showcases detailed comparison videos between TaylorSeer and competing methods on the HunyuanVideo model, which visually demonstrate the significant advantages of our approach. Moreover, TaylorSeer exhibits remarkable versatility by delivering unexpectedly impressive results when extended to super-resolution tasks and audio generation.
| Method | Acceleration | Image | PSNR | SSIM | LPIPS | |
| HiDream | TFLOPs | Speed | Reward | |||
| HiDream-Full | 7780.0 | 1.0 | 1.1285 | - | - | - |
| TeaCache() | 2047.4 | 3.8 | 0.9849 | 28.139 | 0.6036 | 0.565 |
| TaylorSeer() | 1945.0 | 4.0 | 1.0833 | 28.248 | 0.6084 | 0.532 |
12 Performance Analysis at Low Speedup Ratios
Our study reveals that TaylorSeer not only accelerates the inference process but can also surpass the baseline performance, particularly at reduced acceleration ratios. As detailed in Table 1 and Table 2, our empirical results substantiate this finding. Specifically, when applied to FLUX, the ImageReward score improved from 0.9898 to 1.0181. Similarly, for HunyuanVideo, the VBench score increased from 80.66 to 80.74. These results validate the effectiveness of TaylorSeer across the full spectrum of acceleration factors.
We attribute this counter-intuitive performance enhancement to the inherent parameter redundancy within the original large-scale models. This observation is analogous to established techniques such as model pruning and Low-Rank Adaptation (LoRA), where strategically reducing parameter utilization can sometimes lead to improved generalization and overall performance. We hypothesize that TaylorSeer introduces a regularization-like effect during the inference steps. Since both FLUX and HunyuanVideo are based on Flow Matching, the Taylor approximation-based simplification at each step may guide the generation process along a more robust trajectory, effectively regularizing the model’s output. This finding provides valuable insights for both model acceleration and future training methodologies, suggesting that targeted inference-time optimization can unlock latent performance gains.